Samples of machine parts follow a uniform distribution from 35 to 40 grams. What is the standard deviation of this distribution?

  • A :

    1.44 grams

  • B :

    0.42 grams

  • C :

    2.07 grams

  • D :

    3.33 grams
Answer #1

Standard deviation is

\frac{B-A}{\sqrt{12}}

=\frac{40-35}{\sqrt{12}}

=\frac{ 5}{\sqrt{12}}

=1.44 ( A)